OpenStudy (anonymous):

how to find mx+b= y with two points (-2,1) (2,-2)

OpenStudy (anonymous):

find the slope with the slope formula , then use point slope formula with either of the two points and the slope you found.. plug and chug. let me know if you need more explanation..

OpenStudy (anonymous):

(-2-1)/(2-1)=m=slope

OpenStudy (anonymous):

m= -3

OpenStudy (anonymous):

how do you find the b part

OpenStudy (anonymous):

now (y-y1) = m(x-x1) is the point slope formula you know m, and you can use a given point for y1, and x1... I'll use the point (-2,1) y-(1) = (-3)(x-(-2)) y-1 = -3(x+2) y-1 = -3x -6 y = -3x -5

OpenStudy (anonymous):

any time you are given two points... find the slope... then use one of the points with the point slope formula.... Make sure to use parentheses and bring the signs in carefully... hope this helps..